Problem 3: 1-2 Let A = 1-5 1-9 -9 -10 -21 5] 7 14] Find all eigenvalues and corresponding eigenvectors.
Answer #1

the form Let A=(2 - 5 -5 10 T - -21 14 To find eigen valices. Consider (A-dIl = 0 De Determinent zero. -2 -9 57 ſi o -5 to 7=2, -,, & Eigen Vector 5 (A - I] [x] = [0] EzA -9 5 1 [x us -107 T Ly 1-9 -21 14-1 ] -2 -9 -10-2 Rod - 21 Ś n or -5 to 12 -21from ① & ② By Cammers -Y Rule Z - 45 -28+25 48 - 63+60 M eigen Vectors. [x] انا ما ) TON Now for sau [-2+1 -9 5 N 19 -21 14+1for 7 5x i=1 [-2-1 - 7 5 -10- 17 - 21 14-1 to N 71 to di 1-5 -III 1-9 -21 131 -3x -gy + 5z =0 - Soc - lly +72 =0 - -gx-2y +13